OpenStudy (anonymous):

2x + y = 3 x - 2y = -1 If equation one is multiplied by 2 and then the equations are added, the result is _____. 3x = 2 3x = 5 5x = 5

OpenStudy (anonymous):

dont know if u multiply like 2(2x+y=3) or like 2(2)+(2)=3 then add 4x+2y=3+x-2y=-1 or 4+2=3+x-2y=-1

OpenStudy (anonymous):

.................

terenzreignz (terenzreignz):

Multiply like the first @wtbryant 2(2x+y = 3) meaning the entire equation :) and THEN add them.

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

i get 5x+4y=2

OpenStudy (anonymous):

dont know what to do ferom there what i got is not an answer choice

terenzreignz (terenzreignz):

you'll notice that it's a -2y you have in your second equation, not a +2y

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

thanks

terenzreignz (terenzreignz):

:)